OpenStudy (anonymous):

2sin^2x-sinx-1=0, solve for x

OpenStudy (kropot72):

This factorises as follows:\[2\sin ^{2}x-\sin x-1=(2\sin x+1)(\sin x-1)=0\] Can you now solve for x?